sketch the gradient field for the differential equation dy/dt=t-y2 for t=-1......,1

y=-1,.......,1



1
Expert's answer
2021-07-19T18:38:24-0400

Let's calculate the values of the function "f\left(t,y\right)=t-y^2" in the area "A=\left\{\left(t,y\right)\left|-1\le t\le1,\,\,\,-1\le y\le 1\right.\right\}" .To be more detailed, let's take a step of "\Delta t=\Delta y=0.5" . Then,


"t=-1 : \left\{\begin{array}{l}
f(-1,-1)=-1-\left(-1\right)^2=-2\\[0.3cm]
f(-1,-0.5)=-1-\left(-0.5\right)^2=-1.25\\[0.3cm]
f(-1,-0)=-1-0^2=-1\\[0.3cm]
f(-1,0.5)=-1-0.5^2=-1.25\\[0.3cm]
f(-1,1)=-1-1^2=-2
\end{array}\right.\\[0.3cm]
t=-0.5 : \left\{\begin{array}{l}
f(-0.5,-1)=-0.5-\left(-1\right)^2=-1.5\\[0.3cm]
f(-0.5,-0.5)=-0.5-\left(-0.5\right)^2=-0.75\\[0.3cm]
f(-0.5,-0)=-0.5-0^2=-0.5\\[0.3cm]
f(-0.5,0.5)=-0.5-0.5^2=-0.75\\[0.3cm]
f(-0.5,1)=-0.5-1^2=-1.5
\end{array}\right.\\[0.3cm]
t=0 : \left\{\begin{array}{l}
f(0,-1)=0-\left(-1\right)^2=-1\\[0.3cm]
f(0,-0.5)=0-\left(-0.5\right)^2=-0.25\\[0.3cm]
f(0,-0)=0-0^2=-0\\[0.3cm]
f(0,0.5)=0-0.5^2=-0.25\\[0.3cm]
f(0,1)=0-1^2=-1
\end{array}\right.\\[0.3cm]
t=0.5 : \left\{\begin{array}{l}
f(0.5,-1)=0.5-\left(-1\right)^2=-0.5\\[0.3cm]
f(0.5,-0.5)=0.5-\left(-0.5\right)^2=0.25\\[0.3cm]
f(0.5,-0)=0.5-0^2=0.5\\[0.3cm]
f(0.5,0.5)=0.5-0.5^2=0.25\\[0.3cm]
f(0.5,1)=0.5-1^2=-0.5
\end{array}\right.\\[0.3cm]
t=1 : \left\{\begin{array}{l}
f(1,-1)=1-\left(-1\right)^2=0\\[0.3cm]
f(1,-0.5)=1-\left(-0.5\right)^2=0.75\\[0.3cm]
f(1,-0)=1-0^2=1\\[0.3cm]
f(1,0.5)=1-0.5^2=0.75\\[0.3cm]
f(1,1)=1-1^2=0
\end{array}\right."

In the figure it looks like this